Early
Digital Test
Equipment
Transistor Specialties Inc. Apti-Meter 361
Counter
 

Manufacturer: Transistor Specialties Inc.
Model: Apti-Meter 361
Year: 1961
Schematic: RE
 
Max. Rate: 1.3 MHz (measured)
Display: NIXIE
Digits: 6
Logic: discrete transistor Ge-PNP

An instrument from a smaller manufacturer but very well built.

 



- Unit Log -

Serial Number: 398
Year of Manufacture: 1961 (Component date stamp 61??).
Date of Receipt: 2005 Jul 31
State upon Receipt: Fairly clean. MSD NIXIE bad on 5,7,8, but readable. Timebase period incorrect on 10S gate.


Date: 2005 Aug 4
Procedure: Cleaning completed, reassembled.


Date: 2005 Aug 5
Procedure: 180F capacitor for 10S timebase has corroded lead. Replaced with 220F. Trimmer adjusted.


Date: 2005 Aug 6
Procedure: Power-on neon lamp internally covered with metallization. Bulb replaced. 18K limiting resistor replaced with 56K.


Date: 2005 Aug 7
Procedure: V+a filter capacitor (bathtub) and bleeder resistor had been previously replaced, replacements were poorly installed (inadequate support). 8F/220K replaced with 4.7F/390K and moved from location next to crystal oscillator to unused area on power supply board.


Date: 2005 Aug 7
Procedure: Feet had hardened and flattened. Replaced.


Date: 2005 Aug 8
Procedure: Power cord replaced. Hole for strain relief had already been manually trimmed. New lockplate to keep strain relief from turning installed.


Date: 2005 Aug 8
Procedure: Tiny oil drops placed on all control shafts from front.


Date: 2005 Aug 8
Procedure: Fully functional.



- Unit Log -

Serial Number: 399
Year of Manufacture: Component date codes indicate 1961 (6124 on trimmer resistor)
Date of Receipt: 2005 Aug 20
State upon Receipt: Quite clean. Some long scratches on side panels. Unit functional but most NIXIEs have only partial ignition on most numerals. On 10S Test mode, unit only counts to 200.


Date: 2005 Aug 20
Procedure: Cursory cleaning. V+a measures ~ 270V to 310V. The only completely good NIXIE, in digit 10^2, shifted to 10^0.


Date: 2006 Apr 1
Procedure: After an hour or two of being powered on, NIXIEs are much improved.


Date: 2006 Apr 5
Procedure: 180F capacitor for 10S timebase has corroded lead, same as on unit 398. Replaced with 220F. Trimmer adjusted.


Date: 2006 Apr 7
Procedure: Tiny oil drops placed on all control shafts from front.


Date: 2006 Apr 8
Procedure: 18K limiting resistor for power-on neon lamp replaced with 56K.


Date: 2006 Apr 8
Procedure: Crystal oven not warming up. Heater wire open around bend on base at pin 3. New lead installed from pin and heater wire soldered to it. Heater OK.


Date: 2006 Apr 8
Procedure: Feet had hardened and cracked. Replaced.


Date: 2006 Apr 8
Procedure: Cleaning completed, reassembled. Fully functional. NIXIEs are OK after some warm-up.


Date: 2015 Jun 1
Procedure: Powered up, checked. Functional. NIXIE numerals on digits 10^1 to 10^5 still initially only partially illuminate, but are OK after 1/2 to 1 hour warm-up.


Date: 2015 Jun 12
Symptoms: Powered up, snap heard and flash of light, no Nixie illumination.
Procedure: 1N2070 anode-supply rectifier blown up, housing in pieces inside case. Replaced with 1N4007, 300VDC out. Checked first with R in series, filter capacitor seems to be OK.



  .. | Transistor Specialties Inc. Apti-Meter 361 | ..
Early Digital Test Equipment
bhilpert